# Neural Machine Translation of Clinical Text: An Empirical Investigation into Multilingual Pre-Trained Language Models and Transfer-Learning

### 12 Dec 2023 v1

@@ 6 Discussions and Conclusions
To boost the knowledge transformation for digital healthcare and make the most knowledge out of available clinical resources, we explored the state-of-the-art neural langauge models regarding their performances in clinical machine translation.
We investigated a smaller-sized multilingual pre-trained language model (s-MPLM) Marian from Helsinki NLP group, in comparison to two extra-large MPLM (xL-MPLM) NLLB and WMT21fb from Meta-AI. We also investigated transfer-learning possibility in clinical domain translation using xL-MPLM WMT21fb.
We carried out data cleaning and fine-tuning towards clinical domain. We evaluated our work using both automatic evaluation metrics and human expert-based evaluation using HOPE [Gladkoff and Han, 2022] framework.
The experiment also leads to very far-reaching conclusions about MT models and their design, test and applications:
1) The bigger model does not mean that the quality is better, this premise proved to be false, evidently due to the fact that researchers need vast amounts of data to train very large model and very often such data is not clear enough. On the contrary, when we clean the data very well for fine-tuning, we are able to bring the model quality much higher in specific domain, e.g. clinical text. We reached the point when the data quality is more important than the model’s size.
One key take-away for researchers and practitioner from this point is that if it is possible to get 250,000 clean segments in new low resource language, it is capable to fine-tune large language models (LLMs) and get good enough engine in this language. Then, the next step is to continue to get clean data by post-editing translation output from that engine. This is a very important conclusion for “low resource languages”.
2) Automated metrics deliver an illusion of measurement – they are a good tool for iterative stochastic gradient descent during training, but they do not measure quality (only some sort of similarity), are not compatible when any of the underlying factors change, provide results on a non-uniform scale even on their interval of validity, and in general are not sufficiently reliable, and may be misleading. You can’t rely on automatic metrics alone, human translation quality validation is a must, and such validation can deny and reverse the results of automatic measurement.
@@ Acknowledgements
The human evaluation of this work was carried out by Betty Galiano, Marta Martínez Albaladejo, Valeria López Expósito, Carlos Mateos, and Alfredo Madrid; We thanks our human evaluators for their volunteering and hard work. This work has been partially supported by grant EP/V047949/1 “Integrating hospital outpatient letters into the healthcare data space” (funder: UKRI/EPSRC).

### 21 Feb 2024 v2

@@ 6 Discussions and Conclusions
To boost the knowledge transfer for digital healthcare and get the most knowledge out of available clinical resources, we explored the state-of-the-art neural language models regarding their performances in clinical machine translation. We investigated a smaller multilingual pre-trained language model (s-MPLM) Marian from the Helsinki NLP group, in comparison to two extra-large MPLM (xL-MPLM) NLLB and WMT21fb from Meta-AI. We also investigated the transfer-learning possibility in clinical domain translation using xL-MPLM WMT21fb. We carried out data cleaning and fine-tuning in the clinical domain. We evaluated our work using both automatic evaluation metrics and human expert-based evaluation using the HOPE [Gladkoff and Han, 2022] framework.
The experiment has led to some far-reaching conclusions about MT models and their design, test, and applications, in particular:
1) The bigger size of the model does not mean that the quality is better. This premise proved to be false, evidently because researchers need vast amounts of data to train very large models and very often such data is not clear enough. On the contrary, when we clean the data very well for fine-tuning, we can bring the model quality to much higher levels in specific domains, e.g. clinical text. We reached the point where the data quality was more important than the model’s size. One key takeaway for researchers and practitioners from this is that if they can get 250,000 clean segments in a new low-resource language, they can fine-tune large language models (LLMs) and get a good enough engine in this language. Then, the next step is to continue to get clean data by post-editing translation output from that engine. This is a very important implication for “low resource languages.”
2) The automated metrics deliver an illusion of measurement – they are a good tool for iterative stochastic gradient descent during training, but they do not measure quality (only some sort of similarity), are not compatible when any of the underlying factors change, provide results on a non-uniform scale even on their interval of validity, and in general are not sufficiently reliable, and may be misleading. We cannot rely on automatic metrics alone. Instead, human translation quality validation is a must and such validation can deny and reverse the results of automatic measurement.
@@ Acknowledgements
The human evaluation of this work was carried out by Betty Galiano, Marta Martínez Albaladejo, Valeria López Expósito, Carlos Mateos, and Alfredo Madrid. We thank our human evaluators for their volunteering and hard work. We also thank Cristina Sánchez for assisting with double checking the sampled human evaluations. LH and GN are grateful for the support from the grant “Assembling the Data Jigsaw: Powering Robust Research on the Causes, Determinants and Outcomes of MSK Disease.” The project has been funded by the Nuffield Foundation, but the views expressed are those of the authors and not necessarily of the Foundation. Visit www.nuffieldfoundation.org. LH and GN are also supported by the grant “Integrating hospital outpatient letters into the healthcare data space” (EP/V047949/1; funder: UKRI/EPSRC).